Describe your solutions in detail, because partial credit will be given based on your work. If additional information is necessary in solving the problems, DO use your engineering sense and provide reasonable assumptions and explanations. You may use your calculator if necessary. Use a language of your choice. Note: The following information may or may not be used in solving the problems. Gas constant, R-8.314 kPa L/(mol K). 150(1-9)" + 1,75G|, dP The Ergun equation: G (1-4Y where is porosity, p is gas density, dz Dp P9cDp is gas viscosity, G is superficial velocity, and D, is particle diameter. mass
3. (10 pts.) Consider an elementary liquid-phase reaction (A+BC) in a semibatch reactor. The feed stream containing B with a constant concentration (CBo) and a constant flow rate (v0) is slowly fed to a reactor containing pure A with an initial concentration (CAo). Derive three differential equations for the mole balances of A, B, and C. Please provide the steps in detail, because the credits will be given based on the detailed procedure.
Answer #1

Ergvos eqvation (1- )a 1501-dJ4 2 1-5 2 ) ge Dp 3 (ozenay Visco usforce ificant Kosmann vatisn) whan Re< - Gr 1-2150M dz R(1) li4iagea ttion (o exfuo grau fae tos (E) Vo)-Vey0) Reaetini clomtntary Titially Bfeo A Cmd CAD Tmitil CBo Tnitial ContYat2 fox Xeataut B) ol (V=o - D * for b0eloet 4 Ye)v v C)=o - dt YeV 3 - fox EadienD Islo -CCA V- Vd (A (d dt K(A OVEYall TnaybaSimi icyd hy for Yataut B UdBDo- DoCAD t KACBV V= Dot+C SimiLiax y fos c V d Ce - VdCe - k CA CB dt